a.      Using the original values, where K = 100 and L= 25, compute the marginal product of labor at L=25 by calculating how much output would rise if an additional worker were employed.


1
Expert's answer
2021-10-04T10:19:24-0400

Assuming initial quantity produced was 5 and increased to 10 with an additional worker,

Marginal product "=\frac{\Delta L}{\Delta Q}=\frac{26-25}{10-5}=0.2"